# On Continuous Full-Order Integral-Terminal Sliding Mode Control with Unknown A Priori Bound on Uncertainty

### v3

@@ Abstract
This study aims to solve the problem of designing a continuous and finite-time control for a class of nonlinear systems in the presence of matched uncertainty with an unknown apriori bound. First, we propose a Full-Order Integral-Terminal Sliding Manifold (FOITSM) with a conventional (discontinuous) sliding mode to show that it provides the combined attributes of both nonsingular terminal and integral sliding mode algorithms. Secondly, an Adaptive Disturbance Observer (ADO) has been designed to mitigate the impact of the uncertainty. The application of ADO-based Full-Order Integral-Terminal Sliding Mode Control (FOITSMC) substantially reduces chattering in control input under conditionally known matched disturbances. Additionally, the adaptive gains of ADO are updated non-monotonically, avoiding over-bounding of acting disturbances while ensuring global boundedness of state trajectories. Lastly, the proposed algorithm has been effectively applied to the attitude stabilization of a rigid spacecraft.
keywords:
Integral sliding mode control, Terminal sliding mode control, Adaptive disturbance observer, Chattering-free control, Adaptive sliding mode control
journal: ISA Transactions
[…]
[label1]organization=Department of Electrical Engineering, IIT Bombay ,addressline=Powai, city=Mumbai, postcode=400076, state=Maharashtra, country=India
[…]
[label2]organization=Department of Electrical Engineering, IIT Kharagpur ,addressline=Kharagpur, city=Paschim Medinipur, postcode=721302., state=West Bengal, country=India
[…]
[label3]organization=Department of Electrical Engineering, NIT Silchar ,addressline=Silchar, city=Cachar, postcode=788010, state=Assam, country=India
1 Introduction
1.1 Literature survey
@@
Initiation of control amid uncertainties has been one of the main topics of concern for the past few decades. Sliding Mode Control (SMC) [1, 2, 3, 4, 5, 6] is one of the control strategies that converges the state trajectories to a predetermined manifold in finite time. The state trajectories are constrained to evolve in that manifold and thereby, the time evolution of state trajectories becomes insensitive to external disturbances and converges to the origin asymptotically (Conventional sliding mode [7]) or in a finite time (Terminal sliding mode [8, 9, 10, 11, 12, 13, 14]). Manifestation of such control algorithms requires a high frequency switching leading to an undesired phenomenon called ”chattering”. An SMC is generally composed of two phases; the reaching phase, the state trajectories are forced to reach the manifold from its initial condition, and the sliding phase, the state trajectories are compelled to remain invariant to that manifold. In contrast to a conventional linear sliding mode (LSM), the terminal sliding mode control (TSMC) algorithm converges states on finite time and the integral sliding mode (ISM) [15, 16, 17, 18, 19, 20, 21] is devoid of any reaching phase because it ensures the invariance of state trajectories from the initial instant, and hence, guarantees robustness from its inception. Some partial knowledge about the acting external disturbances, preferably the upper bound of the disturbances, must be known in all these algorithms. The apriori knowledge of the matched disturbances is anonymous in several practical applications or bound to differ with time. Our proposed algorithm incorporates this anonymity in the design of the control law, eliminating the reaching phase and significantly reducing chattering in the control input.
@@
Several adaptive sliding mode control (ASMC) algorithms have been proposed to alleviate this barrier. Among many possibilities of adaptation, monotonically increasing the gain [22, 23, 24, 25] associated with the signum structure of the control algorithm may result in high gain [26]. In a recent development [27], a novel ASMC technique is proposed which doesn’t require a known apriori bound on the disturbance and ensures uniformly ultimate boundedness of state trajectories. In [28], an adaptive second-order sliding mode controller has been introduced using the Lyapunov approach, based on the assumption of boundedness of uncertainties rather than the boundedness of their time-derivative. Our proposed methodology further relaxes these assumptions by only requiring the time derivative of acting disturbances to be bounded by an unknown constant value, offering a more flexible and robust solution.
Several approaches are available in the literature regarding continuous adaptive SMC algorithms, preferably adaptive super-twisting control (STC) algorithms [29, 30]. All these approaches can broadly be classified into two main categories. The first category [26, 31, 32, 33] involves increasing the gain of the STC until the second-order sliding (2SM) is obtained; once 2SM has been obtained, adaptive gains stop evolving and thereby, remain constant. The second category [34, 35, 36] comprises the development of adaptive algorithms with minimum gain required to achieve 2SM and, thereby, ensuring that there is no unnecessary overestimation of adaptive gains while ensuring sliding. This category has some advantages over the previous one- specifically from a chattering mitigation viewpoint. To the best of the authors’ knowledge, the number of research articles from a gain minimization perspective is still considerably less and our study aims in this direction using a Lyapunov-based analysis of closed-loop stability to global uniformly ultimately boundedness (GUUB) of state trajectories and estimation errors. We propose a provably robust methodology to manage acting disturbances even when the exact bound of their time-derivative is a constant but unknown. Moreover, this proposed adaptive framework avoids over-estimating adaptive gains. We demonstrate the effectiveness of our proposed control law by comparing our results with an existing methodology and applying it to the attitude stabilization of rigid spacecraft.
1.2 Contributions
In this manuscript, our main contributions are as follows.
  • 1.
    Introducing a Full-order Integral-Terminal Sliding Manifold (FOITSM) deprived of any such reaching phase, ensuring robustness from the initial instant. Besides, in sliding phase, the state trajectories would converge in finite time.
  • 2.
    In the presence of conditionally known disturbances, the control algorithm ensures the globally uniformly boundedness of state trajectories. This conditioning relies on knowledge about the disturbance, specifically that the absolute value of its time-derivative remains bounded. However, the specific bound is unknown to us as an apriori.
  • 3.
    Implementing a disturbance observation technique to mitigate the effect of acting matched disturbances and alleviate the undesired ”chattering” phenomenon.
  • 4.
    Designing adaptive control laws without overestimation of adaptive gains.
1.3 Organisation of the paper
The manuscript unfolds as follows: In Section 2 we delve into the formulation of a layout of a novel FOITSM for a general $n^{\text{th}}$ order chain of integrators afflicted by bounded matched uncertainties with known apriori bounds. An illustrative example is presented in Section 3 for validation of the proposed manifold along with STC. To counterfeit the hurdle of unknown apriori bound on uncertainty, an ADO-based controller has been designed in Section 4. Section 5 shows an illustrative example of the ADO-based FOITSMC control of a 3rd-order chain of integrators. The performance comparison with an existing algorithm is presented in Section 6. In Section 7, an application to attitude Control of a rigid spacecraft has been manifested to show the effectiveness of the proposed disturbance-observer-based controller using the novel manifold in Section 2. Ultimately, the paper draws to a close with a concise summary of conclusions in Section 8.
1.4 Notations
@@
Throughout this paper, $\mathbb{R}$, $\mathbb{R}^{\text{n}}$, $\mathbb{R}^{+}$ and $\mathbb{R}^{\geqslant 0}$ denote, respectively, the set of all real numbers, $n$-dimensional vectors with real components, set of all positive real numbers and nonnegative real numbers. $\mathcal{B}_{r}(x)$ denotes a ball of radius $r$ centered around $x\in\mathbb{R}^{\mathrm{n}}$. The notation $[x]^{\alpha}\text{sgn}(x)$ means $[|x_{1}|^{\alpha}\text{sign}(x_{1}),\cdots,|x_{n}|^{\alpha}\text{sign}(x_{n})]^{T}$, $\forall\alpha\in\mathbb{R},x\in\mathbb{R}^{n}$ and $\text{sgn}(y)$ implies the standard signum function with its argument as $y\in\mathbb{R}$. Further, $|\textbf{x}|=[|x_{1}|,\cdots,|x_{n}|]^{T}$ for $\mathbf{x}\in\mathbb{R}^{n}$ and a summation convention as $a_{1}^{i}\cdots a_{n}^{i}b_{i}^{1}\cdots b_{i}^{m}=\sum_{i}a_{1}^{i}\cdots a_{n}^{i}b_{1}^{i}\cdots b_{m}^{i}$ The function $||\;||:\mathcal{M}\to\mathbb{R}^{+}$ denotes the Euclidean norm on a manifold $\mathcal{M}\subset\mathbb{R}^{\mathrm{n}}$, defined as $||\mathbf{x}||=\sqrt{x^{2}_{1}+x^{2}_{2}+\cdots+x^{2}_{n}}$ for any $\mathbf{x}\in\mathcal{M}$. An open ball of radius $r$ at time $t$ is denoted as $\mathfrak{B}_{r}(\mathbf{x},t)\coloneqq\left\{\mathbf{x}\in\mathcal{M}\;\text{s.t}\;||\mathbf{x}||<r\;\;\text{at time}\;t\right\}$. For the sake of brevity, arguments of a function may be omitted where the context is quite clear.
2 Sliding mode control of nonlinear systems
@@
Consider an $\mathrm{n}^{\mathrm{th}}$ order perturbed chain of integrators which is affine in control
@@
where $\textbf{x}(t)=\left[x_{1}(t),x_{2}(t),\cdots,x_{n}(t)\right]^{T}\in\mathbb{R}^{\mathrm{n}}$ comprises all the states in a vector representation, $f:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the nonlinear drift which is considered to be locally Lipschitz, $b:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the gain for control input, $u(t)\in\mathbb{R}$ is the control input and $d_{0}:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the unknown perturbation representing nonparametric uncertainties and external disturbances. Further
$f(\textbf{x},t)$ $=f_{n}\left(\textbf{x},t\right)+f_{\Delta}\left(\textbf{x},t\right),$
where $f_{n}$ is the nominal component of $f$ and $f_{\Delta}$ is the perturbation from the known nominal value.
@@
The parametric perturbations $f_{\Delta}\left(\textbf{x},t\right)$ is globally bounded, i.e. there exists some real positive constants $f_{\mathrm{max}}$ such that $|f_{\Delta}\left(\textbf{x},t\right)|\leqslant f_{\mathrm{max}}$ where $f_{\mathrm{max}}\in\mathbb{R}^{\geqslant 0}$.
@@
The matched uncertainty $d_{0}\left(\textbf{x},t\right)$ is globally bounded, i.e. $|d_{0}\left(\textbf{x},t\right)|\leqslant d_{\mathrm{max}}$ where $d_{\mathrm{max}}\in\mathbb{R}^{\geqslant 0}$
The primary intent of the sliding mode control approach is to robustly hold on to an effectively chosen constraint, say $s$, through a high-frequency switching control. The purpose is to guide the state trajectories to precisely follow or lie in the vicinity of the chosen constraint.
Definition 2.1 (Ideal sliding [30]).
For $t_{r}\geqslant t_{0}$, if $s(t)\equiv 0,\quad\forall\;t\geqslant t_{r}$, then an ideal sliding has been obtained for system (5) on the sliding manifold $s(t)=0$.
Definition 2.2 (Real sliding [30]).
For $t_{r}\geqslant t_{0}$, $\varepsilon>0$, if $s(t)\in\mathfrak{B}_{\varepsilon}(s,t),\quad\forall\;t\geqslant t_{r}$, then an real sliding has been obtained for system (5) on the sliding manifold $s(t)=0$.
@@
$\left\{\begin{array}[]{l}s=s_{0}(\textbf{x})-z,\\ \dot{z}=-C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})\\ -...-C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}),\end{array}\right.$ (6)
@@
The condition $z(0)$ in (6) ensures that $s=0$ at initial time $t=0$. Hence, the sliding motion occurs from time $t=0$ and thus, there is no reaching phase, unlike conventional TSMC.
Remark 2.
The term “full order” refers to the sliding manifold having a dimension n for an $n^{\mathrm{th}}$ order plant, in contrast to the conventional sliding mode algorithms where the dimension of the sliding manifold is $n-1$ for an $n^{\mathrm{th}}$ order plant.
The dynamical equations (5) and (6) can be combined and represented as
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}\\ \dot{x}_{2}=x_{3}\\ \vdots\\ \dot{x}_{n-1}=s+z\\ \dot{z}=-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})\\ -...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1})\\ \dot{s}=f_{n}(\textbf{x},t)+b(\textbf{x},t)u(\textbf{x})+d(\textbf{x},t)+C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})\\ +C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})+...+C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1})\end{array}\right.$ (15)
where $d\left(\textbf{x},t\right)=d_{0}\left(\textbf{x},t\right)+f_{\Delta}\left(\textbf{x},t\right)$.
Remark 3.
For a specific control $u$, if $s$ is constrained to persistently remain zero for all time $t\geqslant 0$, the dynamical variables in (15) and consequently in (5) follow a trajectory in phase space resembling the solution set of the dynamical equation
$\dot{x}_{n}=$ $-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})$
$-...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1}),$ (16)
which converges to the origin in finite time.
Theorem 2.3.
The state trajectories of the nonlinear system (15) and correspondingly trajectories of (5) under the assumptions 1 and 2, converge to the origin within finite time if the sliding manifold is chosen as in (6). The control input is determined as follows,
$u=$ $b^{-1}(\textbf{x},t)(u_{\text{eqv}}+u_{\text{dis}}),$ (17)
$u_{eqv}=$ $-f_{n}(\textbf{x},t)-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})$
$-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})-...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1}),$ (18)
$u_{dis}=$ $-(\eta+d_{\text{max}}+f_{\text{max}})\text{sgn}(s),$ (19)
where $\eta>0$.
Remark 4.
The discontinuous control $u_{dis}$ in (19) can be replaced by a continuous STC as follows,
$\left\{\begin{array}[]{l}u_{con}=-k_{1}|s|^{\frac{1}{2}}\text{sgn}(s)+v\\ \dot{v}=-k_{2}\text{sgn}(s)\end{array}\right.$ (20)
where $k_{1}$ and $k_{2}$ are parameters defined as $k_{1}=1.5\sqrt{\rho}$ and $k_{2}=1.1\rho$ for $|\dot{d}|\leqslant\rho$.
3 Illustrative example I
To illustrate the fact stated above, let us consider a system of the form,
where the disturbance $d=\text{sin}(2\pi t)$ and $u$ as a control input same as (17), (18) and (20). For simulation, a sliding surface has been considered of the following form.
$s$ $=x_{3}-z,$
$\dot{z}$ $=15\;\text{sgn}(x_{3})|x_{3}|^{\frac{7}{10}}+66\;\text{sgn}(x_{2})|x_{2}|^{\frac{7}{13}}+80\;\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}$ (22)
Following eqs (17)-(19), the corresponding input $u$ is chosen as
$-80\;\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}+u_{con},$
where $u_{con}$ is same as eq (20) with parameters $k_{1}=3.96$ and $k_{2}=7.7$.
(a) Evolution of states on application of STC
(b) Sliding variable of FOITSMC with STC
(c) Input for FOITSMC with STC
Figure 1: 1a and 1b show the convergence of states in finite time and sliding variable remaining $0,\;\forall t\geqslant 0$, hence, demonstrating the effectiveness of FOITSM. On application of STC, a chattering-free control is shown in 1c.
All simulations are done utilising MATLAB R2021a with an ODE 4 (Runge-Kutta) solver assigned a fixed step size of 0.001. Results are illustrated in fig. (1a), (1b) and (1c). In figure (1a), all state trajectories distinctly converge to the origin within finite time. Fig (22) shows the sliding variable starting and persisting at zero, indicating the absence of a reaching phase.. A continuous control $u$, devoid of any singularity as well as chattering (shown in Fig. (1c)), has been applied.
4 Chattering free adaptive SMC with unknown apriori bound on uncertainty
For the plant (21), consider a sliding dynamics $s$ as in eq. (6) with the nature of the disturbance $d\in C^{1}$ as
$|\dot{d}(t)|\leqslant k,$ (23)
where $k$ is a constant but unknown. The choice of $u$ should be such that $s$ becomes zero in a finite time.
Theorem 4.1.
The system (5), characterized by disturbance $d$ as defined in (23), is globally uniformly stable (GUB) and globally uniformly ultimately bounded (GUUB) ([37]) in finite time $t_{r}$

### v4

# Adaptive Disturbance Observer-based Full-Order Integral-Terminal Sliding Mode Control with Unknown A Priori Bound on Uncertainty

@@ Abstract
This study presents a novel, continuous finite-time control strategy for a class of nonlinear systems subject to matched uncertainties with unknown bounds. We propose an Adaptive Disturbance Observer-based Full-order Integral-Terminal Sliding Mode Control (ADO-FOITSMC) to stabilize a chain of integrators in presence of exogenous disturbances whose time derivative is bounded by a constant that is not known a priori. Key features of this approach include a significant reduction in control input chattering and a non-monotonic adaptive law for the observer gains, which prevents overestimation while ensuring the global boundedness of system states. The effectiveness and practical viability of the proposed algorithm are demonstrated through its application to the attitude stabilization of a rigid spacecraft.
I Introduction
@@
Initiation of control in the midst of uncertainties has been one of the main topics of concern for the past few decades. Sliding Mode Control (SMC) (see [1, 2, 3] and the references therein) is one of the control strategies that drives state trajectories to a predefined manifold in finite time (reaching phase) and constrained to evolve in that manifold (sliding phase) which makes the time evolution of state trajectories insensitive to external disturbances. Trajectories converge to the origin asymptotically (Conventional sliding mode [3, 1]) or in a finite time (Terminal sliding mode [4, 5, 6, 7]). However, SMC often involves high-frequency switching, causing undesirable input ”chattering” [1, 2]. Unlike conventional linear sliding mode, TSMC ensures finite-time convergence, while Integral Sliding Mode (ISM) [8, 9, 10, 11, 12] eliminates the reaching phase, ensuring robustness from the outset. These methods typically require prior knowledge of the acting exogenous disturbance, preferably its upper bound, which may be unavailable or time-varying in practice. Our proposed algorithm incorporates this anonymity in the design of the control law, eliminating the reaching phase and significantly reducing chattering in the control input.
@@
Several adaptive sliding mode control (ASMC) algorithms have been developed to address the limitations of traditional SMC. One approach involves monotonically increasing the gain in the signum structure [13, 14, 15, 16], but this can lead to excessively high gains [17]. A novel ASMC technique [18] eliminates the need for a known disturbance bound, ensuring uniformly ultimate boundedness of state trajectories. Another method [19] introduces an adaptive second-order SMC using the Lyapunov approach, assuming bounded uncertainties rather than bounded time derivatives. Our proposed methodology further relaxes these constraints, requiring only that the time derivative of disturbances be bounded by an unknown constant, providing a more flexible and robust solution.
Continuous adaptive super-twisting control (STC) algorithms [20, 21] fall into two categories. The first [17, 22, 23, 24] increases STC gains until second-order sliding mode (2SM) is achieved, then fixes them. The second [25, 26, 27] minimises gains to attain 2SM, reducing chattering and avoiding overestimation. Research on gain minimisation remains limited. Our Lyapunov-based approach ensures global uniform ultimate boundedness (GUUB) of state trajectories and estimation errors, handling disturbances with unknown constant-bounded time derivatives while preventing gain overestimation. We validate its effectiveness through comparisons with existing methods and application to rigid spacecraft attitude stabilisation. In this manuscript, our main contributions are as follows.
  • An Adaptive Disturbance Observer-based Full-order Integral-Terminal Sliding Mode Control (ADO-FOITSMC) that eliminates the reaching phase, ensuring robustness from initiation of control.
  • The proposed algorithm guarantees that the state trajectories are Globally Uniformly Ultimately Bounded (GUUB). This stability is achieved under the relaxed condition that the disturbance’s time derivative is bounded, even when the specific value of this bound is unknown a priori.
  • The proposed adaptive control law avoids gain overestimation while significantly reducing chattering.
@@
Throughout this paper, an open ball of radius $r$ at time $t$ is denoted $\mathfrak{B}_{r}(\mathbf{x},t)=\{\mathbf{x}\in\mathcal{M}:\|\mathbf{x}\|<r\text{ at time }t\}$. For $\alpha\in\mathbb{R}$ and $x\in\mathbb{R}^{n}$, $x]^{\alpha}\text{sgn}(x)=[|x_{1}|^{\alpha}\text{sign}(x_{1}),\dots,|x_{n}|^{\alpha}\text{sign}(x_{n})]^{T}$, with $\text{sgn}(y)$ as the standard signum function for $y\in\mathbb{R}$. The function arguments may be omitted for brevity where the context is clear.
II Problem Formulation
@@
Consider an $\mathrm{n}^{\mathrm{th}}$ order chain of integrators which is affine in control and perturbed by matched disturbance
@@
where $\textbf{x}(t)=\left[x_{1}(t),x_{2}(t),\cdots,x_{n}(t)\right]^{T}\in\mathbb{R}^{\mathrm{n}}$ comprises all the states in a vector representation, $f:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the nonlinear drift which is considered to be locally Lipschitz, $b:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the gain for control input, $u(t)\in\mathbb{R}$ is the control input and $d_{0}:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the unknown perturbation representing nonparametric uncertainties and external disturbances. Further $f(\textbf{x},t)=f_{n}\left(\textbf{x},t\right)+f_{\Delta}\left(\textbf{x},t\right)$, where $f_{n}$ is the nominal component of $f$ and $f_{\Delta}$ is the perturbation from the known nominal value.
@@
The time derivative of parametric perturbation $f_{\Delta}\left(\textbf{x},t\right)$ and matched uncertainty $d_{0}\left(\textbf{x},t\right)$, are assumed to be globally bounded, i.e. there exists some real positive constants $f_{\mathrm{max}}$ and $d_{\mathrm{max}}$ such that $|\dot{f}_{\Delta}\left(\textbf{x},t\right)|\leqslant f_{\mathrm{max}}$ and $|\dot{d}_{0}\left(\textbf{x},t\right)|\leqslant d_{\mathrm{max}}$ where $f_{\mathrm{max}},d_{\mathrm{max}}\in\mathbb{R}^{\geqslant 0}$. For the subsequent analysis, these bounds are consolidated into a single constant $k=f_{\rm max}+d_{\rm max}$.
@@
The specific values of the constant bounds $f_{\mathrm{max}}$ and $d_{\mathrm{max}}$ introduced in Assumption 1 are not known beforehand.
In sliding mode control, the system’s state trajectory is driven towards a predefined sliding manifold, $s(t)=0$. The behavior of the system once it reaches this manifold (or its vicinity) at a finite time $t_{r}$, known as the reaching time, can be categorized as follows.
Definition 1 (Ideal sliding [21]).
An ideal sliding mode represents a scenario where the system’s state trajectory perfectly tracks the sliding manifold after reaching it. Rigorously, an ideal sliding mode is said to be established at time $t_{r}$ if for some finite $t_{r}\geqslant t_{0}$, $s(t)\equiv 0,\;\forall t\geq t_{r}$.
Definition 2 (Real sliding [21]).
For a given, arbitrarily small constant $\varepsilon>0$, a real sliding mode is achieved on the sliding manifold $s(t)=0$ at time $t_{r}\geqslant t_{0}$, if $|s(t)|\leqslant\varepsilon,\;\forall t\geqslant t_{r}$.
@@
$\begin{cases}s=s_{0}(\mathbf{x})-z,\\ \begin{aligned} \dot{z}=&-C_{n}|x_{n}|^{\alpha_{n}}\operatorname{sgn}(x_{n})-C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\operatorname{sgn}(x_{n-1})\\ &-\cdots-C_{1}|x_{1}|^{\alpha_{1}}\operatorname{sgn}(x_{1}),\end{aligned}\end{cases}$ (6)
@@
The condition $z(0)=s_{0}(0)$ ensures $s=0$ at $t=0$, enabling sliding motion from the outset, eliminating the reaching phase typical in conventional TSMC.
Our goal is to synthesize a feedback controller for the system (5) subject to the disturbances as characterised in Assumptions 1 and 2. The goal is to steer the system’s trajectories to, and maintain them within, a bounded vicinity of the sliding manifold (6), thereby realizing a real sliding mode.
III Chattering free adaptive SMC with unknown apriori bound on uncertainty
This section details the design of a feedback control law for system (5) following the assumptions on parametric uncertainties and external disturbances outlined in 1 and 2.
Theorem 1.
The system (5), characterised by disturbance $d$ as assumed in 1 and 2, is globally uniformly stable (GUB) and globally uniformly ultimately bounded (GUUB) ([28]) in finite time $t_{r}$